# Spurious reconstruction from brain activity

### v1

@@ 2.1.1 Observations: Failed replication and convincing reconstruction from random data
These observations raise perplexing questions about the performance and generalizability of recent text-guided reconstruction methods. The significant deterioration in performance when simply changing the dataset from NSD to Deeprecon highlights the need for a deeper understanding of the factors contributing to the success of these methods on the NSD dataset. Moreover, the ability to obtain plausible reconstructions from random brain data by merely generating multiple images and selecting the best ones suggests that there may be fundamental issues with both the evaluation dataset and the components of the reconstruction methods themselves. In the following sections, we will conduct a thorough investigation into the potential problems associated with the NSD dataset and each component of the text-guided reconstruction pipeline.
2.1.2 Lack of diversity in the stimulus set
@@
First, we examined the characteristics and limitations of the NSD dataset itself. To characterize the diversity of stimuli in the datasets and their latent representations, we focused on the CLIP features, which are used as latent features in recent reconstruction methods. We employed uniform manifold approximation and projection (UMAP) (McInnes et al., 2018) to visualize the CLIP text features of the NSD stimuli (see Materials and Methods “UMAP visualization”). The visualization revealed approximately 40 distinct clusters, with considerable overlap between the training and test sets (Figure 4a). Interestingly, we were able to describe the stimulus images in each cluster using a single semantic word, such as airplane, giraffe, or tennis. Despite the NSD containing around 30,000 brain samples per subject, the diversity of the presented stimuli was quite limited to just around 40 semantic concepts. In contrast, the Deeprecon dataset, which was specifically designed to differentiate object categories between training and test data, exhibited less overlap between the two sets (Supplementary Figure S1).
To further investigate the similarity between the training and test stimuli, we conducted an analysis using a state-of-the-art perceptual similarity metric called DreamSim (Fu et al., 2023) (Figure 4b). For each test image in the NSD, we used DreamSim to identify the most similar training images based on their perceptual similarity scores. We observed that the training images extracted using DreamSim were highly similar to the test images in the NSD, not only in terms of semantic concepts but also in terms of overall layout and visual composition. This finding suggests that the NSD test images are heavily biased towards the distribution of the training data, with a significant overlap in the visual and semantic features present in both sets. Such a strong bias raises concerns about the true reconstruction performance of methods evaluated on this dataset, as the impressive results may be largely attributable to the methods’ ability to memorize and reproduce the characteristics of the training data rather than generalizing to novel stimuli.
In contrast, when we applied the same DreamSim-based analysis to the Deeprecon dataset, we found that the extracted training images were substantially different from the test images. This observation indicates a clearer separation between the training and test distributions in the Deeprecon dataset, with the test images containing novel visual and semantic features that are not well-represented in the training set. The distinct differences between the NSD and Deeprecon datasets in terms of stimulus similarity highlight the importance of carefully designing evaluation benchmarks that can effectively assess the generalization capabilities of visual image reconstruction methods.
@@
Figure 4: Dataset analysis. (a) UMAP visualization of CLIP text features in NSD. The center figure shows the scatter plot of the UMAP embedding of CLIP text features in NSD. Green points represent training samples, while orange points represent test samples. The surrounding images show the images in cluster. The presented images in each cluster were randomly selected. (b) Similarity between training and test images using the DreamSim metric. The left side shows results for the NSD, and the right side shows results for the Deeprecon dataset. The top 5 training images with highest similarity to each test image were selected.
2.1.3 Failed zero-shot prediction in the feature space
@@
Given the significant semantic and structural similarities between the NSD test images and the training images, it is crucial to assess whether the predictions made from brain activity truly demonstrate generalization beyond the training set (i.e., zero-shot prediction). To evaluate the zero-shot prediction capability of the translators used in the reconstruction methods, we conducted an $N+1$ way identification analysis, where $N$ represents all of the training samples and $1$ refers to the test sample (Figure 5a). We calculated the percentage of translated features that successfully identified the corresponding test sample in the NSD dataset.
@@
The performance of the CLIP features was nearly $0\%$ (Figure 5b). This extremely low performance suggests that the CLIP feature translator fails to achieve zero-shot prediction, even within the distribution covered by the training set. In other words, the CLIP features translated from brain activity are unable to accurately identify novel test samples, indicating a lack of generalizability beyond the training data. In contrast, the VGG19 features (Simonyan and Zisserman, 2015) employed in the iCNN method demonstrated moderate identification performance at the latent features on the intermediate layers. This finding suggests that the VGG19 features possess a certain degree of generalizability, enabling them to identify test samples that were not encountered during training.
Figure 5: Zero-shot sample identification. (a) Schematic diagram of novel sample identification: calculating the similarity between predicted features and test features or each of the training features, and counting how many predicted features were more similar to the true features than other candidate training features. (b) Results for NSD. The left side shows the identification accuracies evaluated on the CLIP features, which were used in the Brain-diffuser methods. The right side shows the accuracies evaluated on several VGG19 features used in the iCNN method.
@@
Furthermore, we investigated whether the CLIP feature translator enables the prediction of novel semantic concepts not in the training. We redesigned the dataset split to ensure no semantic concepts were shared between them as in the previous zero-shot prediction study (Mitchell et al., 2008; Brouwer and Heeger, 2009). We first applied $k$-means clustering to the UMAP embedding space of the NSD’s CLIP text features (Supplementary Figure S2a). We set the number of clusters as $40$ based on visual inspection of the UMAP results. Based on these clustering results, we performed a hold-out analysis: when predicting samples within a cluster (e.g., the ski cluster), we excluded samples from that cluster in the training set (Figure 6a; hold-out split conditon). As a control, we also prepared a naive data-split condition where the training sample size is the same as in the hold-out split condition but allows overlapping semantic concepts. When we visualized the properties of predicted features in hold-out analysis by transforming the predicted features into the previous UMAP embedding space (Figure 4a), we observed that the predicted features tend to diverge largely from their original clusters and move into other clusters (Figure 6b).
@@
Figure 6: Zero-shot cluster identification. (a) Conceptual diagram of hold-out analysis. For predictions of test samples in a cluster, training samples belonging to the cluster were excluded. The naive split condition serves as a control, using the same number of training samples as in the hold-out split condition but allowing semantic concepts to overlap between training and test sets. (b) Representative example of feature translation under the hold-out split condition. The green and orange points represent training and test samples, respectively. The purple diamonds represent the translated (predicted) latent features. The gray lines represent the transition from the test sample to the corresponding predicted feature. (c) Two types of evaluation of predicted latent features. The left side shows the cluster identification accuracy, which compares predicted features to the average features of training samples within each cluster. The right side shows the pairwise sample identification accuracy, comparing predicted features with the test sample feature or with features of other test samples. (d) Feature prediction results. The four surrounding panels show the semantic clusters observed in the UMAP. In each panel, the upper graphs show the cluster identification performances for CLIP text and CLIP vision features. The expected chance level is $1/40$. The lower graphs show the pairwise sample identification performances for CLIP text and CLIP vision features. The expected chance level is $1/2$.
@@
To quantitatively assess the performance of the feature translator, we employed two metrics: cluster identification accuracy and pairwise sample identification accuracy. Cluster identification accuracy focuses on evaluating the translator’s ability to generate features that correctly identify the semantic cluster to which a test sample belongs (Figure 6c; left). In this analysis, we compare the predicted features of each test sample to the average features of the training samples within each semantic cluster. The accuracy is then calculated as the percentage of predicted features that successfully identify the original semantic cluster of their corresponding test samples in the feature space. Pairwise sample identification accuracy is a commonly used metric in the evaluation of feature prediction and reconstruction performance (Beliy et al., 2019; Shen et al., 2019b, a; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021; Gaziv et al., 2022; Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023; Ferrante et al., 2023; Scotti et al., 2023; Denk et al., 2023; Koide-Majima et al., 2024). This analysis assesses the translator’s ability to generate features that are more similar to the actual test features than to the features of other samples in the test set (Figure 6c; right). For each test sample, we compare its predicted features with two sets of features: the actual test features and the features of another randomly selected sample from the test set. The accuracy is then calculated as the average winning rate of the predicted features against all other test samples, indicating the proportion of cases where the predicted features are more similar to the actual test features than to the features of other samples.
@@
Figure 6d presents a comparison of the identification performance of the feature translator in each semantic cluster between the hold-out and naive split conditions, as measured by two evaluation metrics: cluster identification accuracy and pairwise sample identification accuracy. The cluster identification accuracy in the hold-out split condition exhibited a substantial drop compared to the naive split condition across all semantic clusters. Notably, the cluster identification accuracy was frequently found to be $0\%$ in the hold-out split condition, exposing a severe limitation of the translator in novel semantic concepts that were absent from the training set (see also Supplementary Figure S2b). This finding strongly suggests that the CLIP feature translator functions primarily as a “classifier,” heavily relying on the semantic features learned during the training phase.
@@
Intriguingly, despite the feature translator’s complete failure to identify semantic concepts in the hold-out split condition, the pairwise sample identification accuracy often surpassed the chance level across the semantic clusters (Figure 6d). The insensitivity of the pairwise sample identification accuracy to the translator’s inability to capture semantic concepts can be attributed to its emphasis on comparing the relative similarity between the predicted features and the actual test features, without taking into account the broader categorical context. This can lead to an inflated pairwise sample identification accuracy, potentially misrepresenting the true reconstruction performance and the translator’s ability to generalize to novel clusters. This observation raises concerns about the adequacy of using identification accuracy as the sole metric for evaluating reconstruction performance, a practice that has been widely adopted in many studies (Beliy et al., 2019; Shen et al., 2019a, b; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021; Gaziv et al., 2022; Takagi and Nishimoto, 2023b; Ozcelik and VanRullen, 2023; Ferrante et al., 2023; Scotti et al., 2023; Denk et al., 2023; Koide-Majima et al., 2024).
2.1.4 Failed recovery of a stimulus from its latent features
@@
Finally, we conducted a rigorous evaluation of the generator component, which typically consists of diffusion models in generative AI-based reconstruction methods. To ensure that a visual image reconstruction method has the potential to faithfully reproduce an individual’s perceived visual experiences, it is crucial that the method can recover the original images with a high degree of perceptual similarity when the neural translation from brain activity to latent features is perfect. However, it has been unclear whether generative AI-based reconstruction methods meet this fundamental requirement. To address this question, we performed a recovery check analysis by reconstructing images using the true latent features of target images. Instead of using latent features transformed from brain activity, we directly input the latent features derived from the target images into the generator. While generative AI-based reconstruction methods generated images that were semantically similar to the target, the resulting images were not perceptually similar to the original (Figure 7a). In contrast, iCNN method yielded results that closely resembled the actual target images, demonstrating their superior ability to capture and replicate the original visual content.
Figure 7: Recovery check. (a) Reconstruction from the true latent features. The left column shows the true images, followed by each method’s recovery results. (b) Reconstruction using each component of the Brain-diffuser method. The left column shows the true images, followed by the recovery results of each component of the Brain-diffuser method.
@@
To further investigate the recovery performance, we conducted a recovery check on each latent feature of the Brain-diffuser method (Figure 7b). Interestingly, reconstructions from VDVAE features, which are used for generating initial images in the Brain-diffuser, exhibited a high degree of similarity to the target images. However, the images generated by CLIP features through the diffusion models showed significant deviations from the original targets. These findings suggest that text-guided reconstruction methods may not be well-suited for visual image reconstruction tasks, as they fail to faithfully recover the original visual images. Instead, they tend to create images based on their latent features, such as CLIP features, which can lead to a phenomenon known as “hallucination” in the field of generative AIs. Hallucination refers to an output that appears plausible but is actually incorrect or misleading, raising concerns about the reliability and accuracy of the model (Rawte et al., 2023). These methods seem to prioritize generating semantically similar images rather than faithfully reconstructing the visual content perceived by the individual.
@@
These findings may provide an explanation for why the text-guided reconstruction methods performed well only on the NSD dataset. The results of the case study demonstrated that the text-guided reconstruction methods struggle to reconstruct (Figure 2) or identify (Figure 6) test samples that lie beyond the distribution of the training set. Such limitations suggest that these methods lack true generalization capabilities and are unable to accurately reconstruct novel visual experiences that differ significantly from the examples they were trained on. Moreover, even when the test samples belonged to the same distribution as the training set, the translators of the text-guided reconstruction methods had difficulty correctly identifying those test samples (Figure 5). This observation indicates that the translators may not have learned a sufficiently robust and generalizable mapping between brain activity patterns and the corresponding latent features, further limiting their ability to faithfully reconstruct the perceived visual experiences.
@@
The case study revealed that the NSD test stimuli are highly similar to the training set (Figure 4), with a significant overlap in their visual and semantic features. Given this similarity, the impressive reconstruction results achieved by the recent text-guided reconstruction methods on the NSD dataset should not be interpreted as evidence of zero-shot reconstruction capabilities. Instead, a more plausible interpretation is that these methods primarily function as a combination of “classification” and “hallucination.”
@@
In this context, the translator component of the text-guided reconstruction methods functions primarily as a classifier, predicting categorical information present in the training data rather than capturing the fine-grained details of the visual experience. This limitation may explain why convincing reconstructions can be obtained even from random brain data through post-hoc selection. Due to the limited variety in the outputs generated by the reconstruction models, repeated trials can eventually produce images that are semantically and visually similar to the target stimulus. The apparent plausibility and semantic consistency of these generated images can be attributed to the capabilities of the diffusion model, which learns to generate realistic-looking images based on the semantic information. While these images may seem convincing at first glance, they do not accurately reflect the specific visual experience of the individual. This phenomenon of hallucination raises serious concerns about the reliability and validity of the text-guided reconstruction methods when evaluated on the NSD dataset.
2.2 Formal analysis and simulation
@@
In the above case study, we have identified several issues with these text-guided reconstruction methods and the dataset, especially the cluster structure of CLIP latent features, the lack of diversity in NSD, and the misspecification of the latent representation for image reconstruction, which results in the inability of diffusion models to faithfully recover the original images from their latent features. However, it is crucial to recognize that the findings of the case study are not merely specific to CLIP, NSD, or diffusion models. Instead, these issues likely reflect more fundamental problems that can arise in the development and evaluation of brain decoding and visual image reconstruction methods. Thus in this section, we extend the problems identified in the case study into formal analyses and simulations in generalized settings, aiming to provide a more comprehensive understanding of the factors that contribute to the limitations of current reconstruction methods and explore strategies for mitigating these issues.
2.2.1 Output dimension collapse
@@
First, we revisit the properties of the translator and highlight the issues that arise when the prediction targets have insufficient diversity. Let us consider the general situation of predicting a feature $\mathbf{y}\in\mathbb{R}^{D}$ from a brain activity pattern $\mathbf{x}\in\mathbb{R}^{D}$ using a linear regression model. For the training set, we consider brain activity matrix $X_{\mathrm{tr}}\in\mathbb{R}^{N\times D}$ and feature value matrix $Y_{\mathrm{tr}}\in\mathbb{R}^{N\times D}$, where $X_{\mathrm{tr}}$ consists of $N$ samples of $D$-dimensional brain activity vectors $\mathbf{x}$ and $Y_{\mathrm{tr}}$ consists of $N$ feature value vectors $\mathbf{y}$. We then train the linear (ridge) regression model using this training data. Given a regularization parameter $\lambda$, the weight of ridge regression model is analytically derived as $W=(X_{\mathrm{tr}}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}X_{\mathrm{tr}}^{\top}Y_{\mathrm{tr}}$, where $I$ is the $D\times D$ identity matrix. The predicted feature value $\hat{\mathbf{y}}_{\mathrm{te}}$ for the test brain activity data $\mathbf{x}_{\mathrm{te}}$ can be represented as:
@@
where ${\bf m}=X_{\rm tr}(X_{\rm tr}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}\mathbf{x}_{\mathrm{te}}\in\mathbb{R}^{N\times 1}$ and $\mathbf{y}_{\mathrm{tr}}^{(i)}$ is the $i$th training feature vector. This transformation indicates that the predicted value is always represented as a linear combination of the features in the training set. This property is not limited to ridge regression but is generally applicable to ordinary ridgeless linear regression and related linear models.

### v3

@@ Results
These observations raise perplexing questions about the performance and generalizability of recent text-guided reconstruction methods. The performance significantly deteriorated when we changed from the NSD to Deeprecon dataset. This highlights the need to better understand why these methods succeed with the NSD. Moreover, the ability to obtain plausible reconstructions from random brain data by merely generating multiple images and selecting the best ones suggests that there may be fundamental issues with both the evaluation dataset and the components of the reconstruction methods themselves. In the following sections, we will thoroughly investigate the potential problems associated with the NSD dataset and each component of the text-guided reconstruction pipeline.
Lack of diversity in the stimulus set
@@
First, we examined the characteristics and limitations of the NSD dataset itself. To characterize the diversity of stimuli in the datasets and their latent representations, we focused on the CLIP text features, which are used as latent features in text-guided reconstruction methods. We employed uniform manifold approximation and projection (UMAP) [54] to visualize the CLIP text features of the NSD stimuli (see Materials and Methods “UMAP visualization”). The visualization revealed approximately $40$ distinct clusters, with considerable overlap between the training and test sets (Fig 4A). Interestingly, we were able to describe the stimulus images in each cluster using a single semantic word, such as airplane, giraffe, or tennis. Despite the NSD containing around $30,000$ brain samples per subject, the diversity of the presented stimuli was quite limited to just around $40$ semantic concepts. Note that we performed UMAP visualization using the parameters recommended in the official guide for clustering, but a similar cluster structure was also observed with the default parameters (S1 Fig). In contrast, the Deeprecon dataset, which was specifically designed to differentiate object categories between training and test data, exhibited less overlap between the two sets (S2 Fig).
@@
Fig 4: Dataset diversity and similarity between training and test stimuli. (A) UMAP visualization of CLIP text features in the NSD. The center figure shows the scatter plot of the UMAP embedding of CLIP text features. The gray points represent training samples, while the orange points represent test samples. The surrounding images were randomly selected from each cluster. (B, C) Similarity between training and test images. For each example test image (the one on the left of each row), the five training images with the highest similarity were selected using the DreamSim metric and displayed. This analysis was performed for the NSD (B) and Deeprecon (C) datasets, using the same procedure.
To further investigate the similarity between the training and test stimuli, we analyzed using a state-of-the-art perceptual similarity metric called DreamSim [55]. For each test image in the NSD, we used DreamSim to identify the most similar training images based on their perceptual similarity scores. We found that the training images extracted using DreamSim were highly similar to the test images in the NSD, not only in terms of semantic concepts but also in terms of overall layout and visual composition (Fig 4B). These findings suggest that the distribution of NSD test images is heavily biased towards that of the training images, with a significant overlap in the visual and semantic features present in both sets. Such a strong bias raises concerns about the actual reconstruction performance of methods evaluated on this dataset. The impressive results may be largely attributable to the methods’ ability to memorize and reproduce the characteristics of the training stimuli rather than generalizing them to novel stimuli.
In contrast, when we applied the same DreamSim-based analysis to the Deeprecon dataset, we found that the extracted training images were substantially different from the test images (Fig 4C). This observation supports the careful design of the train-test split in the Deeprecon dataset, ensuring that the test images contain novel visual and semantic features that are not well-represented in the training set. The distinct differences between the NSD and Deeprecon datasets in terms of stimulus similarity highlight the importance of carefully designing evaluation benchmarks that can effectively assess the generalization capabilities of visual image reconstruction methods.
Failed zero-shot prediction in the feature space
@@
Given the significant semantic and structural similarities between the NSD test images and the training images, it is crucial to assess whether the predictions derived from brain activity truly demonstrate generalization beyond the training set (i.e., zero-shot prediction). To evaluate the zero-shot prediction capability of the translators used in the reconstruction methods, we conducted an $(N+1)$-way identification analysis, where $N$ represents all of the training samples ($8,859$ samples for subject 1 in the NSD) and $1$ refers to the test sample (Fig 5A). Specifically, we calculated the similarity (correlation) between the predicted (translated) features and the true features of corresponding test samples, as well as the similarities between the predicted features and the true features corresponding to each of the N training samples. We then examined whether the predicted features correctly identified (showed the highest similarity to) the test true features. We calculated the percentage of predicted features that successfully identified the corresponding test sample in the NSD dataset.
Fig 5: Evaluation of zero-shot sample identification performance. (A) Schematic diagram of novel sample identification. Similarity is calculated between predicted (translated) features and test features, as well as between predicted features and each of the training features. Identification accuracy is determined by how often the predicted features show the highest similarity to the true test features compared to all training features. (B) Zero-shot identification performance using predicted CLIP features (text and vision) from the Brain-Diffuser method, evaluated on the NSD dataset. (C) Zero-shot identification performance using predicted features from several intermediate layers of the VGG19 model employed in the iCNN method, also evaluated on the NSD dataset. For both (B) and (C), the identification task involved $8,859$ training samples plus one test sample, resulting in an $8,860$-way identification. The chance performance level is thus $1/8860$.
@@
The performance of the CLIP features was nearly 00% (Fig 5B). The CLIP features predicted from brain activity were unable to accurately identify novel test samples, even within the distribution covered by the training set. This extremely low performance suggests that the CLIP feature translator fails to achieve zero-shot prediction. This finding raises questions about the effectiveness of CLIP features for true zero-shot prediction tasks in brain decoding. In contrast, the VGG19 features [56] employed in the iCNN method demonstrated moderate identification performance at the latent features on the intermediate layers (Fig 5C). This finding suggests that the VGG19 features possess a certain degree of generalizability, enabling them to identify test samples not encountered during training.
@@
Furthermore, we investigated whether the CLIP feature translator enables the prediction of novel semantic concepts not in the training. We redesigned the dataset split to ensure no semantic concepts were shared between them as in the previous zero-shot prediction study [17, 18]. We first applied $k$-means clustering to the UMAP embedding space of the NSD’s CLIP text features (S3 FigA). We set the number of clusters as $40$ based on visual inspection of the UMAP results. Based on these clustering results, we performed a hold-out analysis: when predicting samples within a cluster (e.g., the ski cluster), we excluded samples from that cluster in the training set (Fig 6A right; hold-out split condition). As a control, we also prepared a naive data-split condition where the training sample size is the same as in the hold-out split condition but allows overlapping semantic concepts (Fig 6A left; naive split condition). When we visualized the properties of predicted features in hold-out analysis by transforming them into the previous UMAP embedding space (Fig 4A), we observed that the predicted features tend to diverge largely from their original clusters and move into other clusters (Fig 6B).
@@
Fig 6: Cluster hold-out analysis. (A) Hold-out procedure. When predicting a test sample from a specific cluster, all training samples from that cluster are excluded. The naive split condition serves as a control, using the same number of training samples but allowing cluster overlap between training and test sets. (B) Examples of predictions. Under the hold-out split condition, gray and dark orange points represent training and test samples, respectively. Light orange points indicate translated (predicted) latent features. Black lines connect true features to their corresponding predictions. (C) Cluster identification procedure. The predicted features are compared with the average features of each cluster, and the most similar cluster is selected. (D) Cluster identification results. Surrounding panels display CLIP text and vision feature performance in the hold-out and naive split conditions of the four representative semantic clusters. The chance level is $1/40$. (E) Pairwise sample identification procedure. Predicted features are compared with the true test sample and another randomly chosen test sample. A sample more similar to the prediction is selected. (F) Pairwise sample identification results. The accuracy means the average winning rate of the true samples against all of the other test samples. Results are presented similarly to cluster identification, with a chance level of $1/2$.
@@
To quantitatively assess the performance of the feature translator, we employed two identification metrics. The first metric is cluster identification accuracy. Cluster identification accuracy focuses on evaluating the translator’s ability to generate features that correctly identify the semantic cluster to which a test sample belongs (Fig 6C). In this analysis, we compare the predicted features of each test sample to the average features of the training samples within each semantic cluster. The accuracy is then calculated as the percentage of predicted features that successfully identify the original semantic cluster of their corresponding test samples in the latent feature space.
@@
The cluster identification accuracy in the hold-out split condition exhibited a substantial drop compared to the naive split condition across all semantic clusters (Fig 6D). Notably, the cluster identification accuracy was frequently 00% in the hold-out split condition (see also S3 Fig.B for all cluster results). These results expose a severe limitation of the translator when dealing with novel semantic concepts absent from the training set. This suggests that the CLIP feature translator primarily functions as a “classifier;” its prediction (translation) heavily relies on predefined semantic features used in a training phase rather than generalizing to new concepts.
The second metric is pairwise sample identification accuracy. Pairwise sample identification accuracy is a commonly used metric in the evaluation of feature prediction and reconstruction performance [39, 29, 34, 36, 40, 41, 44, 45, 46, 57, 58]. This analysis assesses the translator’s ability to predict target features that are more similar to the actual test features than to the features of other samples in the test set (Fig 6E). For each test sample, we compare its predicted features with two sets of features: the actual test features and the features of another randomly selected sample from the test set. The accuracy is then calculated as the average winning rate of the predicted features against all other test samples, indicating the proportion of cases where the predicted features are more similar to the actual test features than those of other samples.
@@
Intriguingly, despite the feature translator’s complete failure to identify semantic concepts in the hold-out split condition, the pairwise sample identification accuracy often surpassed the chance level across the semantic clusters (Fig 6F). The insensitivity of the pairwise sample identification accuracy to the translator’s inability to capture semantic concepts can be explained by its focus on relative similarity. Specifically, the similarity compares the predicted features to the actual test features, without taking into account the broader general context. This can lead to an inflated pairwise sample identification accuracy, potentially misrepresenting the actual reconstruction performance and the translator’s ability to generalize to novel clusters. This observation raises concerns about the adequacy of using identification accuracy as the sole metric for evaluating reconstruction performance, a practice that has been widely adopted in many studies.
Failed recovery of a stimulus from its latent features
@@
Finally, we conducted a rigorous evaluation of the generator component, which typically consists of diffusion models in text-guided reconstruction methods. To ensure that a visual image reconstruction method has the potential to faithfully reproduce an individual’s perceived visual experiences, it is crucial that the method can recover the original images with a high degree of perceptual similarity when the neural translation from brain activity to latent features is perfect. However, it has been unclear whether recent text-guided reconstruction methods meet this fundamental requirement. To address this question, we performed a recovery check analysis by reconstructing images using the true latent features of target images. Instead of using latent features translated from brain activity, we directly input the latent features derived from the target images into the generator.
Text-guided methods produced images semantically similar to targets but not perceptually similar (Fig 7A), while the iCNN method yielded results that closely resembled the actual target images. These findings suggest that text-guided reconstruction methods may prioritize semantic similarity over perceptual accuracy. In contrast, the iCNN method appears to have a superior ability to capture and replicate original visual content, indicating potential advantages in preserving fine-grained visual details.
Fig 7: Recovery check. (A) Reconstruction from the true latent features. The leftmost column displays the original images. Subsequent columns show reconstruction results from the StableDiffusionReconstruction, Brain Diffuser, and iCNN methods. (B) Component-wise reconstruction of the Brain-Diffuser method. The leftmost column shows the original images. The following columns present reconstruction results using individual components of the Brain-Diffuser method (VDVAE, CLIP text, and CLIP vision features). Each row represents a different test image. The reconstruction results reveal the upper bound of reconstruction performance for each method and highlight potential limitations in the latent feature representations or generative processes.
@@
To investigate the recovery performance further, we conducted a recovery check on each latent feature of the Brain-Diffuser method (Fig 7B). Interestingly, reconstructions from VDVAE features, which are used for generating initial images in the Brain-Diffuser, exhibited a high degree of similarity to the target images. However, the images generated by CLIP features through the diffusion models showed significant deviations from the original targets. These findings suggest that text-guided reconstruction methods may not be well-suited for visual image reconstruction tasks, as they fail to faithfully recover the original visual images. Instead, they tend to create images based on their latent features, such as CLIP features, which can lead to a phenomenon known as “hallucination” in the field of generative AIs. Hallucination refers to an output that appears plausible but is actually incorrect or misleading, raising concerns about the reliability and accuracy of the model [59]. Text-guided reconstruction methods seem to prioritize generating semantically similar images rather than faithfully reconstructing the visual content perceived by the individual.
@@
The above findings may provide an explanation for why the text-guided reconstruction methods performed well only on the NSD dataset (Fig 2A). The results of the case study demonstrated that the text-guided reconstruction methods struggle to reconstruct (Fig 2B) or identify (Fig 6) test samples that lie beyond the distribution of the training set. Such limitations suggest that these methods lack true generalization capabilities and are unable to accurately reconstruct novel visual experiences that differ significantly from the examples they were trained on. Moreover, even when the test samples belonged to the same distribution as the training set, the translators of the text-guided reconstruction methods had difficulty correctly identifying those test samples (Fig 5). This observation indicates that the translators may not have learned a sufficiently robust and generalizable mapping between brain activity patterns and the corresponding latent features, further limiting their ability to faithfully reconstruct the perceived visual experiences.
@@
This case study also revealed that the NSD test stimuli are highly similar to the training set, with a significant overlap in their visual and semantic features (Fig 4). Given this similarity, the impressive reconstruction results achieved by the recent text-guided reconstruction methods on the NSD dataset should not be interpreted as evidence of zero-shot reconstruction capabilities. Instead, a more plausible interpretation is that these methods primarily function as a combination of “classification” and “hallucination.”
@@
In this context, the translator component of the text-guided reconstruction methods functions primarily as a classifier, predicting categorical semantic information present in the training phase rather than capturing the fine-grained details of the visual experience. This limitation may explain why convincing reconstructions can be obtained even from random brain data through post-hoc selection (Fig 3B). Due to the limited variety in the outputs generated by the reconstruction models, repeated trials and post-hoc selection can eventually find images that are semantically and visually similar to the target stimulus. The apparent plausibility and semantic consistency of these generated images can be attributed to the capabilities of the diffusion model, which learns to generate realistic-looking images based on semantic information. While these images may seem convincing at first glance, they do not accurately reflect the specific visual experience of the individual. This phenomenon of hallucination raises serious concerns about the reliability and validity of the text-guided reconstruction methods when evaluated on the NSD dataset. Although we evaluated only two reconstruction methods in this case study, it is important to recognize that any reconstruction methods evaluated only by NSD [46, 48, 60] can potentially have similar classification and hallucination problems.
Formal analysis and simulation
@@
Building on the issues identified in our case study, we examine these challenges in a more general context. We have identified several issues with these text-guided reconstruction methods and the dataset, including the cluster structure of CLIP latent features, the lack of diversity in NSD, and the misspecification of the latent representation for image reconstruction. These issues resulted in the inability of diffusion models to faithfully recover the original images from their latent features. However, it is crucial to recognize that the findings of the case study are not merely specific to CLIP, NSD, or diffusion models. Instead, these issues likely reflect more fundamental problems that can arise in the development and evaluation of brain decoding and visual image reconstruction methods. Thus in this section, we extend the problems identified in the case study into formal analyses and simulations in generalized settings, aiming to provide a more comprehensive understanding of the factors that contribute to the limitations of current reconstruction methods and explore strategies for mitigating these issues.
Output dimension collapse
Multivariate linear regression models are widely used in constructing decoding and encoding models of the brain. These models are often regarded as capable of independent and compositional predictions for each target, as they create separate regression models for individual targets without sharing the weights between them. However, this expectation is not generally true. This is particularly evident when input variables are shared [33, 35, 34, 44, 45]. In this section, we demonstrate this assertion by examining the problem of predicting multiple targets using linear regression models with shared inputs.
@@
Let us consider predicting a feature vector $\mathbf{y}\in\mathbb{R}^{D}$ from a brain activity pattern $\mathbf{x}\in\mathbb{R}^{D}$ using a linear regression model. For the training set, we consider brain activity matrix $X_{\mathrm{tr}}\in\mathbb{R}^{N\times D}$ and feature value matrix $Y_{\mathrm{tr}}\in\mathbb{R}^{N\times D}$, where $X_{\mathrm{tr}}$ consists of $N$ samples of $D$-dimensional brain activity vectors $\mathbf{x}$ and $Y_{\mathrm{tr}}$ consists of $N$ feature vectors $\mathbf{y}$. We then train a linear (ridge) regression model using this training data. Given a regularization parameter $\lambda$, the weight of the ridge regression model is analytically derived as $W=(X_{\mathrm{tr}}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}X_{\mathrm{tr}}^{\top}Y_{\mathrm{tr}}$ where $I$ is the $D\times D$ identity matrix. The predicted feature vector $\hat{\mathbf{y}}_{\mathrm{te}}$ for the test brain activity data $\mathbf{x}_{\mathrm{te}}$ can be represented as:
@@
where ${\bf m}=X_{\rm tr}(X_{\rm tr}^{\top}X_{\mathrm{tr}}+\lambda I)^{-1}\mathbf{x}_{\mathrm{te}}\in\mathbb{R}^{N}$, $m_{i}$ is the $i$th element of $\mathbf{m}$, and $\mathbf{y}_{\mathrm{tr}}^{(i)}$ is the $i$th training feature vector. This transformation indicates that the predicted value is always represented as a linear combination of the target features in the training set. This property is not limited to ridge regression but generally applies to ordinary ridgeless linear regression and related linear models.